Calculate price elasticity of demand using point elasticity method when the company is in equilibrium and interpret the result.
1
Expert's answer
2019-05-14T09:00:38-0400

Point elasticity is the price elasticity of demand at a specific point on the demand curve instead of over a range of it.

If Qd = a - b×P, then Ed = -b×P/Q.